OpenStudy (anonymous):

from a committee of 3 girls and 6 boys, two are chosen at randonm to lead the committee.what is rthe probability that both people chosen will be boys

OpenStudy (anonymous):

There are 6+3=9 people altogether, so the probability of the first person being a boy is 6/9=2/3, because there are 6 boys out of 9 people. The probability of the second person being a boy is now 5/8, because there are only 5 boys left, out of 8 people, as one boy has been taken out. So, \[\left(\begin{matrix}2 \\ 3\end{matrix}\right) \times \left(\begin{matrix}5 \\ 8\end{matrix}\right) =\left(\begin{matrix}10 \\ 24\end{matrix}\right) = \left(\begin{matrix}5 \\ 12\end{matrix}\right)\]

OpenStudy (anonymous):

I toss a coin and roll a number Find P(heads and even number)

OpenStudy (anonymous):

luke has 3 pair of trouser and 3 pairs of boots. How many different outfits can he wear

OpenStudy (anonymous):

I roll a standard number cube Find P (number greater than 1)

OpenStudy (anonymous):

P(heads) =1/2 P(even number)=3/6=1/2 \[\left(\begin{matrix}1 \\ 2\end{matrix}\right) \times \left(\begin{matrix}1 \\ 2\end{matrix}\right) = \left(\begin{matrix}1 \\ 4\end{matrix}\right)\]

OpenStudy (anonymous):

Say each pair of trousers was a, b and c, and each pair of boots was 1, 2 and 3. The possibilities are: a1 a2 a3 b1 b2 b3 c1 c2 c3 That's 9 different outfits.

OpenStudy (anonymous):

I'm guessing a 'standard number cube' is a dice, so: P(>1)=5/6

OpenStudy (anonymous):

a bag contains 7 yellow, 6 blue, and 3 red marbles. 2 marbles are drawn random, one at a time, without replacement. which is the probability that both are blue

OpenStudy (anonymous):

a bag contains seven blue, 8 red, and 4 white marbles. find the probability of picking 3 red marbles if each marble is returned to the bag before the next marble picked

OpenStudy (anonymous):

the probability that Shaina is on time for school is 1/2 find the probability that Shaina arrives on time for school for the 5 days. express the answer as percent